13.In a field there are twice as many birds as cows and one-third as many horses as birds.There are 640 legs in all.If no animal is missing a leg, how many animals are there?
a.60
b.120
c.220
d.260
e.300

Answer by mathmandotcom(31)

Let x= number of cows
then 2x=the number of birds
and(1/3)2x=the number of horses
if each cow and horse has 4 legs and each bird has 2 legs then
total number of legs=4(number of cows+number of horses)+2(number of birds)=640
#legs=4(x+2/3*x)+2(2x)=640 multiply and simplify we get
640=4x+8/3*x+4x
640=8x+8/3*x multiply by 3
1920=24x+8x
1920=32x divide by 32 we get
x=60 # of cows
2(60)=120= # of birds
1/3(2*60)=40= # of horses
cows +birds+horses=60+120+40=220
total number of animals is 220 answer is C
-----------